OpenStudy (anonymous):

all calculations leading to an answeA student collected the following data for a fixed volume of gas: Temperature (⁰C) Pressure (mm of Hg) 10 726 20 750 40 800 70 880 100 960 150 ??? Fill in the missing data point. Show r. help please

OpenStudy (anonymous):

The General Law of Gases states that P1*V1/T1 = P2*V2/T2. As Volume is constant, we can write this as P1/T1 = P2/T2, P1 and T1 are the pressure and temperature at time 1. P2 and T2 are the pressure and temperature at time 2. You can take any pair of values as P1 and T1 and find P2 = P1T2/T1, once you sum 273.15 to each temperature to get kelvins. P1 = 726 mm Hg T1 = 10 + 273.15 = 283.15 K; T2 = 150 + 273.15 = 423.15 K. Then, P2 = 726 * 423.15/283.15 = 1085.96 mm Hg 10,726 ) , (100, 960) slope is (960 - 726) / (100 - 10) = 2.6 y = 2.6 x + b find y intercept 726 = 2.6*10 + b 726 - 26 = b 700 = b y = 2.6x + 700 this means in celsius , pressure = 2.6 * temperature + 700 now plug in x = 150 y = 2.6(150) + 700 = 1090 I got so far

OpenStudy (anonymous):

but it said my answer was wrong

OpenStudy (anonymous):

try to get im working on it im kinda struggling..